# Estimating Semantic Alphabet Size for LLM Uncertainty Quantification

### v1

@@ Abstract
Many black-box techniques for quantifying the uncertainty of large language models (LLMs) rely on repeated LLM sampling, which can be computationally expensive. Therefore, practical applicability demands reliable estimation from few samples. Semantic entropy (SE) is a popular sample-based uncertainty estimator with a discrete formulation attractive for the black-box setting. Recent extensions of semantic entropy exhibit improved LLM hallucination detection, but do so with less interpretable methods that admit additional hyperparameters. For this reason, we revisit the canonical discrete semantic entropy estimator, finding that it underestimates the “true” semantic entropy, as expected from theory. We propose a modified semantic alphabet size estimator, and illustrate that using it to adjust discrete semantic entropy for sample coverage results in more accurate semantic entropy estimation in our setting of interest. Furthermore, our proposed alphabet size estimator flags incorrect LLM responses as well or better than recent top-performing approaches, with the added benefit of remaining highly interpretable.
@@ 1 Introduction
We provide empirical evidence that canonical discrete semantic entropy underestimates the “true” SE for typical sample sizes, on average (Figure 2). To address this limitation, we first suggest a straightforward modification to existing alphabet size estimators, where we draw, in part, from population ecology. Then, we use it to adjust entropy estimates for sample coverage, resulting in reduced bias (Figure 2) and more accurate estimation of the “true” semantic entropy, compared to other black-box semantic entropy estimators (Table 1).
Finally, we evaluate our suggested estimators for inaccuracy classification in sentence-length LLM responses. In our experiments, the aforementioned coverage-adjusted estimator outperforms other explicit discrete estimators of semantic entropy. More strikingly, the alphabet size estimator alone outranks all other considered UQ methods in overall strength - except for one, from which we do not statistically distinguish performance, after accounting for uncertainty in establishing an overall rank of methods. Our results indicate that semantic alphabet size estimation, which is highly interpretable, can perform as well or better than state-of-the art methods in LLM inaccuracy detection.
@@ 2 Preliminaries
Semantic entropy (SE), introduced by Kuhn et al. (2023), aims to quantify intrinsic LLM uncertainty with a three-step procedure:
@@
Because the distribution over semantic sets is not known, it is approximated from response probabilities $p(d|q,\theta)$ using so-called “Rao–Blackwellized Monte Carlo integration”:
@@
Recent work is said to have generalized semantic entropy. Two pertinent examples are Kernel Language Entropy (KLE) and Semantic Nearest Neighbor Entropy (SNNE), which reported state-of-the-art performance for incorrectness classification (Nikitin et al., 2024; Nguyen et al., 2025). These stronger estimators may come at the expense of interpretability, in part owing to implementation complexity (e.g., applying a kernel embed graph nodes in KLE) and introduction of additional hyperparameters and design choices (e.g., similarity function and scale parameter in SNNE). We elaborate on KLE and SNNE definitions and implementation details in Appendix A.5.
@@ 3 Methods
The total number of semantic equivalence classes (i.e., the semantic alphabet size $|S|$) that may be ellicited from an LLM in response to a prompt is generally unknown, and not all categories are necessarily observed in the sample (i.e., $k<|S|$). For instance, under a simple Zipfian model of semantic category-abundance, we expect at least one category to be unobserved for sample size $n=10$ if $|S|>4$ (Appendix B). This situation is known as the under-sampled regime, where the empirical distribution over semantic categories can be less surprising than than the true one. In other words, the plugin method for DSE may underestimate LLM uncertainty, which we illustrate in Figure 2. For this reason, we are interested in methods for estimating semantic alphabet size and adjusting for it when estimating semantic entropy.
@@
$\hat{\mathbb{H}}_{CS}=-\sum_{i=1}^{k}\frac{\hat{C}_{GT}\hat{p}_{i}\log(\hat{C}_{GT}\hat{p}_{i})}{1-(1-(\hat{C}_{GT}\hat{p}_{i}))^{n}}.$ (8)
This so-called “Chao-Shen” estimator is consistent and less biased than many empirical alternatives (Vu et al., 2007; Pinchas et al., 2024).
@@ 4.1 Experimental settings
Following Farquhar et al. (2024), we focus on fine-tuned models - in our case, the instruction-tuned models of Gemma-2-9B (Team et al., 2024), Llama-3.1-8B (Grattafiori et al., 2024), Mistral-v0.3-7B, and Phi-3.5-3.8B (Abdin et al., 2024). We perform text generation at two temperatures, for distinct purposes. Following related prior work, we sample at temperature $\tau=1.0$ to calculate uncertainty scores and again at $\tau=0.1$ to obtain a “best guess” response for assessing the correctness of LLM responses (Farquhar et al., 2024; Nikitin et al., 2024; Nguyen et al., 2025).
Figure 2: Illustrating underestimation in discrete semantic entropy calculation with typical sample sizes. The ratios of semantic entropy estimators with varying sample size ($n=5,10,25,50,75,100$) to white-box semantic entropy with $n=100$ (denoted $SE^{*}$) are shown, with values below 11 suggesting underestimation (dotted grey line). The estimators displayed are the plugin estimator of canonical discrete semantic entropy (i.e., Equation 5, dotted indigo line) and the “hybrid” semantic entropy estimator of Equation 10 (solid indigo line). Results are averaged across queries within each dataset, then uniformly averaged across datasets. Log scale is used on the x-axis to highlight differences between estimators with smaller sample sizes. Instances with only one observed semantic category at $n=100$, resulting in a denominator of 0, are ignored.
@@
We consider question-answering datasets wherein each query has either one or multiple correct answer(s). For the former, we use the validation sets of HotpotQA (Yang et al., 2018) and SQuAD 2.0 (Rajpurkar et al., 2018; 2016), which contain $1852$ and $11864$ exemplars, respectively. Our experiments require extensive sampling of LLM responses, and BEC demands combinatorially many language model calls. For these reasons, and because SQuAD 2.0 is much larger than the other data sets, our experiments using it are performed on a random $20\%$ subset. We also consider queries with a wider variety of possible correct answers. The first such dataset is BioASQ, a biomedical QA benchmark with $4719$ exemplars. Each question in BioASQ has between $1$ and $15$ reference answers. We also prepare a small ($131$ exemplars) supplementary dataset called Plethora Of accepTable cATegOries (POTATO), for which the number of possible correct semantic categories has an even larger range (up to $722$). We discuss POTATO in greater detail in Section A.2.
@@ 4.2 Metrics and baselines
For a given prompt-LLM pair, $|S|$ and the true probability distribution over semantic equivalence classes are unknown, so semantic alphabet size and semantic entropy have no ground truths. Instead, we assume that white-box semantic entropy with a large number of samples ($n=100$), denoted $SE^{*}$, is interchangeable with the true estimand (i.e., $SE^{*}$ well-represents the true entropy over all possible semantic categories). When illustrating the degree to which $SE^{*}$ is underestimated (i.e., Figure 2), we report the ratio of the estimator of interest to $SE^{*}$. Since information entropy is non-negative, values below $1$ indicate underestimation. Our primary evaluation metric for entropy estimators, however, is mean-squared error (MSE): we assess the MSE between an estimate using the small sample size ($n=10$) and $SE^{*}$.
@@
We also assess the ability of estimators to classify LLM responses as “correct” or “incorrect,” sometimes referred to as “hallucination” or “confabulation” detection (Farquhar et al., 2024). We calculate the area under the receiver operating characterstic curve (AUROC), corresponding to the probability that a randomly selected incorrect LLM response has a higher uncertainty score than a randomly selected correct response. Empirical AUROC values are point estimates, which themselves have uncertainty. Nikitin et al. (2024) point out that such uncertainty is strongly driven by the model and dataset, rather than UQ method, motivating head-to-head comparisons and evaluation by win rate. We advance a similar approach, but we rely on Bradley-Terry latent strength scores (Zermelo, 1929; Bradley & Terry, 1952), calculated via minorization-maximization (Caron & Doucet, 2012; Hunter, 2004), allowing us to establish an overall rank of methods.
We also employ a Monte Carlo procedure that attempts to account for uncertainty in estimating both AUROC and strength scores. First, we fit Gaussian uncertainty distributions about AUROC point estimates from the $95\%$ confidence intervals (CIs) obtained via DeLong’s method (DeLong et al., 1988; Sun & Xu, 2014), a modeling assumption we justify by the approximate normality of the U statistic (Mann & Whitney, 1947). [^1] In principle, samples from these distributions may exist outside the $[0,1]$ range, but this is unlikely in our case, given the AUROC CI bounds (Figure 7). For each model-dataset pair, we simulate $L=100$ matches between each pair of methods by comparing AUROC values sampled from their corresponding uncertainty distributions. After obtaining strength scores from the simulated matches, we calculate CIs about the Bradley-Terry scores and establish $95\%$ CIs about the ranks of latent strengths using the methods of Gao et al. (2023). We provide additional details in Appendix A.6.
@@
Because our priority is black-box uncertainty estimation, we compare with other black-box approaches. For semantic entropy estimation, we compare our presented hybrid approach (i.e., Equation 10) to the other explicit discrete semantic entropy estimators: the canonical discrete approach of Farquhar et al. (2024) (i.e., the plugin estimator, Equation 5) and the coverage-adjusted estimator of Chao & Shen (2003) (i.e., Equation 8).
For incorrectness classification, we consider the three aforementioned explicit semantic entropy estimators, four alphabet size estimators, and three other uncertainty methods. The alphabet size estimators are the number of semantic categories (i.e., NumSets) (Kuhn et al., 2023), the alphabet size estimator converted from the Good-Turing sample coverage estimator (Good, 1953), the spectral estimator of Lin et al. (2024) (i.e., $U_{EigV}$), and our presented hybrid alphabet size estimator (i.e., Equation 9). The other uncertainty methods are Predictive Entropy (PE, see Appendix A.5) (Kadavath et al., 2022), SNNE (Nguyen et al., 2025), and KLE (Nikitin et al., 2024).
@@ 4.3 Results
Dataset Estimator Gemma-2-9B Llama-3.1-8B Mistral-7B Phi-3.5-3.8B
HotpotQA $\widehat{\mathbb{H}}_{Plugin}$ 0.47 ± 0.02 0.72 ± 0.03 0.60 ± 0.03 0.62 ± 0.03
$\widehat{\mathbb{H}}_{CS}$ 0.39 ± 0.02 0.57 ± 0.03 0.46 ± 0.03 0.47 ± 0.03
$\widehat{\mathbb{H}}_{Hybrid}$ 0.30 ± 0.02 0.45 ± 0.02 0.39 ± 0.02 0.39 ± 0.02
SQuAD 2.0 $\widehat{\mathbb{H}}_{Plugin}$ 0.69 ± 0.03 0.84 ± 0.03 1.40 ± 0.04 1.46 ± 0.04
$\widehat{\mathbb{H}}_{CS}$ 0.51 ± 0.03 0.60 ± 0.03 0.86 ± 0.06 0.91 ± 0.06
$\widehat{\mathbb{H}}_{Hybrid}$ 0.43 ± 0.02 0.50 ± 0.02 0.68 ± 0.03 0.74 ± 0.03
POTATO $\widehat{\mathbb{H}}_{Plugin}$ 0.42 ± 0.08 0.71 ± 0.12 1.75 ± 0.15 1.96 ± 0.16
$\widehat{\mathbb{H}}_{CS}$ 0.30 ± 0.07 0.52 ± 0.11 0.94 ± 0.29 1.57 ± 0.55
$\widehat{\mathbb{H}}_{Hybrid}$ 0.27 ± 0.06 0.42 ± 0.09 0.70 ± 0.11 0.72 ± 0.11
BioASQ $\widehat{\mathbb{H}}_{Plugin}$ 1.66 ± 0.03 1.81 ± 0.03 2.06 ± 0.03 1.85 ± 0.02
$\widehat{\mathbb{H}}_{CS}$ 0.96 ± 0.04 1.05 ± 0.05 1.31 ± 0.08 0.98 ± 0.04
$\widehat{\mathbb{H}}_{Hybrid}$ 0.78 ± 0.02 0.78 ± 0.02 0.72 ± 0.02 0.82 ± 0.02
Table 1: Empirically evaluating the accuracy of explicit discrete semantic entropy estimators. Values reflect MSE between the estimated value using $n=10$ samples and white-box semantic entropy with $n=100$ (i.e., $SE^{*}$). The lowest MSE value for each model-dataset pair is shown in bold. Intervals, shown in grey, reflect $95\%$ CIs via the standard error of the mean.
Canonical DSE underestimates “true” semantic entropy.
The plugin estimator for information entropy $\hat{\mathbb{H}}_{plugin}$ has a theoretically-established negative bias (Basharin, 1959; Harris, 1975), which largely governs its mean-squared error (Antos & Kontoyiannis, 2001). We are interested in observing this phenomenon empirically, where we do not have access to the true distribution over semantic equivalence classes. In Figure 2, we compare plugin DSE estimates to $SE^{*}$ (dotted lines), indicating that the canonical plugin DSE approach underestimates its quantity of interest for practical sample sizes. Such underestimation can be problematic for reliable UQ, because drawing a large number of samples from an LLM can be costly and time-consuming at scale.
@@
Figure 3: Heatmap illustrating the portion of model-dataset pairs, rounded to the nearest integer, for which a row’s method achieved a larger AUROC point estimate than a column’s method. Uncertainty methods are organized into three groups: (i) semantic alphabet size estimators, (ii) semantic entropy estimators, and (iii) other uncertainty methods. The hybrid discrete semantic entropy estimator consistently outperforms other DSE estimators, and the hybrid semantic alphabet size estimator consistently outperforms other alphabet size estimators.
@@
In Figure 2, we observed that $\hat{\mathbb{H}}_{Hybrid}$ underestimated its target quantity by less than canonical plugin DSE for a range of sample sizes. Detailed results, broken out by dataset, are shown in Appendix C (Figure 6), where similar patterns are exhibited. We perform a more granular comparison of semantic entropy estimator accuracy with $n=10$ in Table 1, where $\hat{\mathbb{H}}_{Hybrid}$ consistently achieves the lowest MSE among explicit discrete semantic entropy estimators across four models and four datasets.
We improve upon prior work on incorrectness detection.
@@
LLM uncertainty estimation is often employed as a proxy for hallucination, confabulation, or incorrectness detection among LLM responses (Kuhn et al., 2023; Farquhar et al., 2024). Following Nikitin et al. (2024), we calculate pairwise AUROC win rates between models – for each model-dataset pair, a win is recorded in method $i$’s favor over method $j$ if the corresponding AUROC for method $i$ is greater than that of method $j$ (Figure 3). We characterize overall performance of each uncertainty method via Bradley-Terry latent strength scores. Figure 4A displays maximum likelihood estimates of the methods’ strength scores from pairwise AUROC comparisons. Here, $\widehat{|S|}_{Hybrid}$ achieves the highest strength score, followed closely by KLE.
Figure 4: Establishing overall performance of ten UQ methods on incorrectness detection. (A) Bradley-Terry latent strength scores from pairwise comparison of AUROC point estimates. (B) Bradley-Terry latent strength scores after accounting for uncertainty in estimating AUROC. Error bars are “conservative” CIs about strength scores, which may be slightly stricter than $95\%$; see Section A.6 for details. (C) For each method, we establish $95\%$ CIs about the rank of Bradley-Terry latent strength MLEs (Gao et al., 2023) for the incorrectness detection task; see Section A.6 for details. We highlight i) semantic alphabet size estimators, ii) explicit discrete semantic entropy estimators, and iii) other uncertainty estimators. The interval [a, b] denotes all integers from $a$ to $b$, inclusively. The CI for SNNE reflects [6, 6] and is extended for readability.
We highlight the uncertainty in both AUROC and Bradley-Terry estimation in Figure 4B (CIs about Bradley-Terry estimates, which may be slightly stricter than $95\%$) and Figure 4C ($95\%$ CIs around each method’s rank by Bradley-Terry score). We observe that the hybrid estimators presented herein outperform estimators of the same type - i.e., $\widehat{|S|}_{Hybrid}$ achieves the highest strength score among semantic alphabet size estimators, and $\widehat{|S|}_{Hybrid}$ does the same among explicit semantic entropy estimators. Finally, we find that $\widehat{|S|}_{Hybrid}$ shares the [1, 2] CI with KLE (Figure 4C), a conceptually complex approach that may be more difficult to interpret. We do not statistically distinguish the incorrectness detection performance of $\widehat{|S|}_{Hybrid}$ from that of KLE, with both methods holding top performance, after accounting for uncertainty in our rank-generating procedure. Our results are consistent with Kuhn et al. (2023)’s observation that the number of observed semantic categories is itself “a reasonable uncertainty measure.”

### v2

@@ Abstract
Many black-box techniques for quantifying the uncertainty of large language models (LLMs) rely on repeated LLM sampling, which can be computationally expensive. Therefore, practical applicability demands reliable estimation from few samples. Semantic entropy (SE) is a popular sample-based uncertainty estimator with a discrete formulation attractive for the black-box setting. Recent extensions of SE exhibit improved LLM hallucination detection, but do so with less interpretable methods that admit additional hyperparameters. For this reason, we revisit the canonical discrete semantic entropy (DSE) estimator, finding that it underestimates the “true” semantic entropy, as expected from theory. We propose a modified semantic alphabet size estimator, and illustrate that using it to adjust DSE for sample coverage results in more accurate SE estimation in our setting of interest. Furthermore, we find that two semantic alphabet size estimators, including our proposed, flag incorrect LLM responses as well or better than many top-performing alternatives, with the added benefit of remaining highly interpretable.
@@ 1 Introduction
We provide empirical evidence that canonical discrete semantic entropy (DSE) underestimates the “true” semantic entropy (SE) for typical sample sizes, on average (Figure 2). To address this limitation, we first suggest a straightforward modification to existing alphabet size estimators, where we draw, in part, from population ecology. Then, we use it to adjust entropy estimates for sample coverage, resulting in reduced bias (Figure 2) and more accurate estimation of the “true” SE, compared to other black-box SE estimators (Table 1).
Finally, we evaluate UQ estimators for inaccuracy classification in sentence-length LLM responses. In our experiments, the aforementioned coverage-adjusted estimator outperforms other explicit estimators of SE (Figure 3). More strikingly, two alphabet size estimators outrank all other black-box UQ methods considered but one, after accounting for uncertainty in establishing an overall rank of methods. Our results indicate that semantic alphabet size estimation, which is highly interpretable, can perform competitively with many state-of-the art methods in LLM inaccuracy detection.
@@ 2 Preliminaries
SE, introduced by Kuhn et al. (2023), aims to quantify intrinsic LLM uncertainty with a three-step procedure:
@@
Because the distribution over semantic sets is not known, it is approximated from response probabilities $p(d|q,\theta)$ using so-called “Rao–Blackwellized Monte Carlo integration” (RBMCI):
@@
Recent work is said to have generalized SE or DSE. Two pertinent examples are Kernel Language Entropy (KLE) and Semantic Nearest Neighbor Entropy (SNNE), which reported state-of-the-art performance for incorrectness classification (Nikitin et al., 2024; Nguyen et al., 2025). These stronger estimators may come at the expense of interpretability, in part owing to implementation complexity (e.g., applying a kernel to embed graph nodes in KLE) and introduction of additional hyperparameters and design choices (e.g., similarity function and scale parameter in SNNE). We elaborate on KLE and SNNE definitions and implementation details in Appendix A.5.
@@ 3 Methods
The total number of semantic equivalence classes (i.e., the semantic alphabet size $|S|$) that may be ellicited from an LLM in response to a prompt is generally unknown, and not all categories are necessarily observed in the sample (i.e., $k<|S|$). For instance, under a simple Zipfian model of semantic category-abundance, we expect at least one category to be unobserved for sample size $n=10$ if $|S|>4$ (Appendix B). This situation is known as the under-sampled regime, where the empirical distribution over semantic categories can be less surprising than than the true one. In other words, the plugin method for DSE may underestimate LLM uncertainty, which we illustrate in Figure 2. For this reason, we are interested in methods for estimating semantic alphabet size and adjusting for it when estimating SE.
@@
$\hat{\mathbb{H}}_{CS}=-\sum_{i=1}^{k}\frac{\hat{C}_{GT}\hat{p}_{i}\log(\hat{C}_{GT}\hat{p}_{i})}{1-(1-(\hat{C}_{GT}\hat{p}_{i}))^{n}},$ (8)
[…]
@@ 4.1 Experimental settings
Following Farquhar et al. (2024), we focus on fine-tuned models - in our case, the instruction-tuned models of Gemma-2-9B (Team et al., 2024), Gemma-3-12B (Team et al., 2025), Llama-3.1-8B (Grattafiori et al., 2024), Mistral-v0.3-7B, and Phi-3.5-3.8B (Abdin et al., 2024). We perform text generation at two temperatures, for distinct purposes. Following related prior work, we sample at temperature $\tau=1.0$ to calculate uncertainty scores and again at $\tau=0.1$ to obtain a “best guess” response for assessing the correctness of LLM responses (Farquhar et al., 2024; Nikitin et al., 2024; Nguyen et al., 2025). With the exception of Figure 2, which iterates over several sample sizes, the results shown in the main body of this work use a sample size of $n=10$, which is also found in prior work on SE (Kuhn et al., 2023; Farquhar et al., 2024; Nikitin et al., 2024; Nguyen et al., 2025). Due to its reduced computational burden compared to larger sample sizes and prevalence in the relevant literature, we will refer to $n=10$ as a “practical” or “typical” sample size.
Figure 2: Illustrating underestimation in discrete semantic entropy (DSE) calculation with typical sample sizes. The ratios of DSE estimators with varying sample size ($n=5,10,25,50,75,100$) to white-box SE with $n=100$ (denoted $SE^{*}$) are shown, with values below 11 suggesting underestimation (dotted grey line). The estimators displayed are the plugin estimator of canonical DSE (i.e., Equation 5, dotted indigo line) and the “hybrid” DSE estimator of Equation 10 (solid indigo line). Results are averaged across queries within each dataset, then uniformly averaged across datasets. Log scale is used on the x-axis to highlight differences between estimators with smaller sample sizes. Instances with a denominator of 0 are ignored.
@@
We consider question-answering datasets wherein each query has either one or multiple correct answer(s). For the former, we use the validation sets of HotpotQA (Yang et al., 2018) and SQuAD 2.0 (Rajpurkar et al., 2018; 2016), which contain $7405$ and $11864$ exemplars, respectively. Our experiments require extensive sampling of LLM responses, and BEC demands combinatorially many language model calls. For these reasons, and because SQuAD 2.0 is much larger than the other data sets, our experiments using it are performed on a random $20\%$ subset. HotpotQA has substantial scale, as well, due to its inclusion of contextual information supplementing the queries. For estimator evaluation (Figure 2 and Table 1), we use the first $25\%$ of HotpotQA’s queries; for incorrectness detection (Figure 3), we use the full dataset. We also consider queries with a wider variety of possible correct answers. The first such dataset is BioASQ, a biomedical QA benchmark with $4719$ exemplars. Each question in BioASQ has between $1$ and $15$ reference answers. We also prepare a small ($131$ exemplars) supplementary dataset called Plethora Of accepTable cATegOries (POTATO), for which the number of possible correct semantic categories has an even larger range (up to $722$). We discuss POTATO in detail in Section A.2.
@@ 4.2 Metrics and baselines
For a given prompt-LLM pair, $|S|$ and the true probability distribution over semantic equivalence classes are unknown, so semantic alphabet size and SE have no ground truths. Instead, we assume that white-box SE with a large number of samples ($n=100$), denoted $SE^{*}$, is interchangeable with the true estimand (i.e., $SE^{*}$ well-represents the true entropy over all possible semantic categories). We elaborate on this choice in Appendix Section C. When illustrating the degree to which $SE^{*}$ is underestimated (i.e., Figure 2), we report the ratio of the estimator of interest to $SE^{*}$. Since information entropy is non-negative, values below $1$ indicate underestimation. Our primary evaluation metric for entropy estimators, however, is mean-squared error (MSE): we assess the MSE between an estimate using the small sample size ($n=10$) and $SE^{*}$.
@@
We also assess the ability of estimators to classify LLM responses as “correct” or “incorrect,” sometimes referred to as “hallucination” or “confabulation” detection (Farquhar et al., 2024). We calculate the area under the receiver operating characteristic curve (AUROC), corresponding to the probability that a randomly selected incorrect LLM response has a higher uncertainty score than a randomly selected correct response. Empirical AUROC values are point estimates, which themselves have uncertainty. Nikitin et al. (2024) point out that such uncertainty is strongly driven by the model and dataset, rather than UQ method, motivating head-to-head comparisons and evaluation by win rate. We advance a similar approach, but we rely on Bradley-Terry latent strength scores (Zermelo, 1929; Bradley & Terry, 1952), calculated via minorization-maximization (Caron & Doucet, 2012; Hunter, 2004), allowing us to establish an overall rank of methods. We also employ a Monte Carlo procedure that attempts to account for uncertainty in estimating both AUROC and strength scores. First, we fit Gaussian uncertainty distributions about AUROC point estimates from the $95\%$ confidence intervals (CIs) obtained via DeLong’s method (DeLong et al., 1988; Sun & Xu, 2014), a modeling assumption we justify by the approximate normality of the U statistic (Mann & Whitney, 1947). [^1] In principle, samples from these distributions may exist outside the $[0,1]$ range, but this is unlikely in our case, given the AUROC CI bounds (Figure 10). For each model-dataset pair, we simulate $L=100$ matches between each pair of methods by comparing AUROC values sampled from their corresponding uncertainty distributions. After obtaining strength scores from the simulated matches, we calculate CIs about the Bradley-Terry scores and establish $95\%$ CIs about the ranks of latent strengths using the methods of Gao et al. (2023). We provide additional details in Appendix A.6.
@@
Because our priority is black-box uncertainty estimation, we compare with other black-box methods. For SE estimation, we compare our hybrid approach (Equation 10) to the other explicit black-box SE estimators: the canonical discrete approach of Farquhar et al. (2024) (the plugin estimator, Equation 5) and the coverage-adjusted estimator of Chao & Shen (2003) (Equation 8). For incorrectness classification, we consider the three aforementioned explicit black-box SE estimators, four alphabet size estimators, and four other uncertainty methods. The alphabet size estimators are the number of semantic categories (i.e., NumSets) (Kuhn et al., 2023), the alphabet size estimator converted from the Good-Turing sample coverage estimator (Good, 1953), the spectral estimator of Lin et al. (2024) (i.e., $U_{EigV}$), and our hybrid alphabet size estimator (i.e., Equation 9). The other uncertainty methods are Predictive Entropy (PE, see Appendix A.5) (Kadavath et al., 2022), SNNE (Nguyen et al., 2025), KLE (Nikitin et al., 2024), and white-box SE (Kuhn et al., 2023).
@@ 4.3 Results
Dataset Estimator Gemma-2 (9B) Gemma-3 (12B) Llama-3.1 (8B) Mistral-v0.3 (7B) Phi-3.5 (3.8B)
HotpotQA $\widehat{\mathbb{H}}_{Plugin}$ 0.46 ± 0.02 0.09 ± 0.01 0.68 ± 0.03 0.59 ± 0.03 0.61 ± 0.03
$\widehat{\mathbb{H}}_{CS-GT}$ 0.39 ± 0.02 0.08 ± 0.01 0.56 ± 0.03 0.46 ± 0.03 0.47 ± 0.03
$\widehat{\mathbb{H}}_{Hybrid}$ 0.30 ± 0.02 0.06 ± 0.01 0.45 ± 0.02 0.39 ± 0.02 0.39 ± 0.02
SQuAD 2.0 $\widehat{\mathbb{H}}_{Plugin}$ 0.68 ± 0.03 0.18 ± 0.02 0.78 ± 0.03 1.37 ± 0.04 1.43 ± 0.04
$\widehat{\mathbb{H}}_{CS-GT}$ 0.51 ± 0.03 0.15 ± 0.01 0.60 ± 0.03 0.86 ± 0.06 0.91 ± 0.06
$\widehat{\mathbb{H}}_{Hybrid}$ 0.43 ± 0.02 0.12 ± 0.01 0.50 ± 0.02 0.68 ± 0.03 0.75 ± 0.03
POTATO $\widehat{\mathbb{H}}_{Plugin}$ 0.39 ± 0.08 0.35 ± 0.08 0.62 ± 0.11 1.71 ± 0.15 1.91 ± 0.15
$\widehat{\mathbb{H}}_{CS-GT}$ 0.30 ± 0.07 0.29 ± 0.07 0.51 ± 0.11 0.94 ± 0.29 1.58 ± 0.55
$\widehat{\mathbb{H}}_{Hybrid}$ 0.27 ± 0.06 0.24 ± 0.06 0.42 ± 0.09 0.71 ± 0.11 0.73 ± 0.11
BioASQ $\widehat{\mathbb{H}}_{Plugin}$ 1.64 ± 0.03 1.79 ± 0.03 1.68 ± 0.03 2.04 ± 0.03 1.82 ± 0.02
$\widehat{\mathbb{H}}_{CS-GT}$ 0.96 ± 0.04 1.10 ± 0.07 1.06 ± 0.05 1.31 ± 0.08 0.99 ± 0.04
$\widehat{\mathbb{H}}_{Hybrid}$ 0.78 ± 0.02 0.69 ± 0.02 0.80 ± 0.02 0.73 ± 0.02 0.83 ± 0.02
Table 1: Empirically evaluating the accuracy of explicit discrete semantic entropy (DSE) estimators. Values reflect MSE between the estimated value using $n=10$ samples and white-box semantic entropy with $n=100$ (i.e., $SE^{*}$). The lowest MSE value for each model-dataset pair is shown in bold. Intervals, shown in grey, reflect $95\%$ CIs via the standard error of the mean.
Figure 3: Establishing overall performance of ten UQ methods on incorrectness detection. (A) Bradley-Terry latent strength scores from pairwise comparison of AUROC point estimates. (B) Bradley-Terry latent strength scores after accounting for uncertainty in estimating AUROC. Error bars are “conservative” CIs about strength scores, which may be slightly stricter than $95\%$; see Section A.6 for details. (C) For each method, we establish $95\%$ CIs about the rank of Bradley-Terry latent strength MLEs (Gao et al., 2023) for the incorrectness detection task; see Section A.6 for details. We highlight (i) semantic alphabet size estimators, (ii) black-box discrete semantic entropy (DSE) estimators, and (iii) other uncertainty estimators, which include white-box (PE, SE) and black-box methods (SNNE, KLE). The interval [a, b] denotes all integers from $a$ to $b$, inclusively. The CI for SNNE reflects [6, 6] and is extended for readability.
Canonical discrete semantic entropy underestimates “true” semantic entropy.
The plugin estimator for information entropy $\hat{\mathbb{H}}_{plugin}$ has a theoretically-established negative bias (Basharin, 1959; Harris, 1975), which largely governs its mean-squared error (Antos & Kontoyiannis, 2001). We are interested in observing this phenomenon empirically, where we do not have access to the true distribution over semantic equivalence classes. In Figure 2, we compare plugin DSE estimates to $SE^{*}$, indicating that the canonical plugin DSE approach underestimates its quantity of interest for practical sample sizes. Such underestimation can be problematic for reliable UQ, because drawing a large number of samples from an LLM can be costly and time-consuming at scale.
@@
In Figure 2, we observe that, on average, $\hat{\mathbb{H}}_{Hybrid}$ misestimates its target quantity by less than canonical plugin DSE for a range of sample sizes. We perform a more granular comparison of SE estimator accuracy with $n=10$ in Table 1, where $\hat{\mathbb{H}}_{Hybrid}$ consistently achieves the lowest MSE among explicit discrete SE estimators across five models and four datasets. We also find that $\hat{\mathbb{H}}_{Hybrid}$ slightly over-shoots $SE^{*}$ on average for the 12B-parameter Gemma model. Detailed results, broken out by dataset in Appendix Figure 7, reveal that this overshooting is limited to the POTATO dataset. In fact, this appears to be driven by just three anomalous instances, examined in Appendix D. Excluding instances with near-zero “true” entropy (i.e., $SE^{*}<0.005$) recovers the pattern exhibited by the other models (Figure 9).
Our proposed semantic entropy estimator outperforms other explicit semantic entropy estimators at incorrectness detection.
@@
LLM uncertainty estimation is often employed as a proxy for hallucination, confabulation, or incorrectness detection among LLM responses (Kuhn et al., 2023; Farquhar et al., 2024). Following Nikitin et al. (2024), we calculate pairwise AUROC win rates between models – for each model-dataset pair, a win is recorded in method $m_{i}$’s favor over method $m_{j}$ if the corresponding AUROC for $m_{i}$ is greater than that of $m_{j}$ (Figure 4). We characterize overall performance of each method via Bradley-Terry latent strength scores. Figure 3A displays maximum likelihood estimates of the methods’ strength scores from pairwise AUROC comparisons. We highlight the uncertainty in both AUROC and Bradley-Terry estimation in Figure 3B (CIs about Bradley-Terry estimates, which may be slightly stricter than $95\%$) and Figure 3C ($95\%$ CIs around each method’s rank by Bradley-Terry score). Here, $\widehat{|H|}_{Hybrid}$ (Equation 10) achieves the highest rank among explicit estimators of semantic entropy, including white-box SE.
Alphabet size estimators alone can outperform entropy estimators at incorrectness detection.
After accounting for uncertainty in our rank-generating procedure, two semantic alphabet size estimators - namely, $\widehat{|S|}_{Hybrid}$ and $U_{EigV}$ - outperform many more complex black-box uncertainty methods. Three methods have $95\%$ rank CIs that include the top position: alphabet size estimators $\widehat{|S|}_{Hybrid}$ (rank CI of [1, 4]) and $U_{EigV}$ (rank CI of [1, 3], which attains the largest latent strength estimate), as well as KLE (rank CI of [1, 3]) (Figure 3C), with all three methods outperforming white-box SE and other explicit SE estimators overall. Our results recall Kuhn et al. (2023)’s observation that the number of observed semantic categories is itself “a reasonable uncertainty measure.”